OpenStudy (anonymous):

How can I find X here? |Solving for X| 2x + 1 = x2 + 4x + 18

OpenStudy (anonymous):

@tHe_FiZiCx99

OpenStudy (anonymous):

@Zarkon

OpenStudy (anonymous):

Ok, looks like I can use the quadratic formula. Can someone help me apply it?

OpenStudy (anonymous):

@jigglypuff314

jigglypuff314 (jigglypuff314):

first get all the stuff to one side

OpenStudy (anonymous):

Do you mean combine like terms?

OpenStudy (anonymous):

And get X to one side?

jigglypuff314 (jigglypuff314):

subtract 2x and 1 from both sides then quadratic

OpenStudy (anonymous):

Ok, thanks. Check my answer when I'm done?

jigglypuff314 (jigglypuff314):

sure :)

OpenStudy (anonymous):

Ok, I have the problem set up.

OpenStudy (anonymous):

\[x = \frac{ -2x plusorminus \sqrt{4x^2 - 4(x^2 * 17)} }{ 2x^2 }\]

OpenStudy (anonymous):

plusorminus = ±

OpenStudy (anonymous):

Ugh.

OpenStudy (anonymous):

Did I set up the equation correctly?

jigglypuff314 (jigglypuff314):

\[\pm \]

OpenStudy (anonymous):

Yes, I couldn't find it lol

jigglypuff314 (jigglypuff314):

ummm the quadratic equation is just for the coefficients the whole point was to take the x's out of there to solve for it :)

jigglypuff314 (jigglypuff314):

\[\frac{ -2\pm \sqrt{(2)^2-4(1)(17)} }{ 2(1) }\]

jigglypuff314 (jigglypuff314):

oh wait O_O there are no real answers!!!

jigglypuff314 (jigglypuff314):

@superdude123 ... are you learning imaginary?

OpenStudy (anonymous):

Sorry!

OpenStudy (anonymous):

Why are there no real answers?

OpenStudy (anonymous):

I graphed it an they don't intersect, which means there is no answer.

jigglypuff314 (jigglypuff314):

square root ( 4 - (4*17) ) = squareroot(negative) = imaginary number

OpenStudy (anonymous):

So should I say that the answer is extraneous?

OpenStudy (anonymous):

Or the equation, rather.

jigglypuff314 (jigglypuff314):

um more like there are no real solutions

OpenStudy (anonymous):

Prime?

jigglypuff314 (jigglypuff314):

umm I just think it's no real solutions unless you have options?

OpenStudy (anonymous):

Nope, that's fine. Thanks for your help. :)